OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -2x-11y=5.

OpenStudy (freethinker):

slope intercept form" y = mx + b your m in the mx is the slope your b is the y-intercept

OpenStudy (anonymous):

oh, nvm. I know

OpenStudy (anonymous):

I got y=2/11x+5/11 Is that correct?

OpenStudy (freethinker):

so if you have a problem that looks like -4x - 22y = 5 you need to make sure it is in the form of y = mx + b in this case, you will have to isolate -22y first by adding 4x into both sides of the equation -4x + (4x) - 22y = 5 + (4x) that leaves you with -22y = 4x +5 we are looking for y so it would mean both sides has to be divided by -22 leaving us with y = -(4/22) x -5/22 you can simplify it further into y = (2/11) x - 5/22

OpenStudy (freethinker):

it is not correct @Emily778